OpenStudy (anonymous):

check my answer please? .A GE lightbulb is supposed to last for 1,200 hours. In fact, light bulbs of this type last only 1,185 hours with a standard deviation of 70 hours. What is the probability that a sample of 100 lightbulbs will have an average life of at least 1,200 hours? 1 in 2.41

OpenStudy (lilshane):

i think your right

OpenStudy (jordan123321):

yeah

OpenStudy (anonymous):

Thank you

OpenStudy (perl):

first find the probability that an individual lightbulb has a life of at least 1200 hours

OpenStudy (anonymous):

@perl do we know the lifetime of a light bulb is normally distributed though?

OpenStudy (perl):

good point :) , it doesnt say

OpenStudy (anonymous):

We can probably assume it in this case, but it doesn't hurt to be careful

OpenStudy (anonymous):

so do you agree with my answer?

jimthompson5910 (jim_thompson5910):

you need to use the xbar distribution for this one

OpenStudy (perl):

yes because the sample is size 100

jimthompson5910 (jim_thompson5910):

the xbar distribution is the distribution of sample means. You are looking for the area under the curve to the right of x = 1200. This curve is normally distributed with mean = 1185, standard deviation = sigma/sqrt(n) = 70/sqrt(100) = 7

OpenStudy (anonymous):

0.214?

OpenStudy (perl):

$$ \Large { P( \bar X \geq 1200) = P( Z_\bar{X} \geq \frac{1200 - 1185}{\frac{70}{\sqrt{100}}} ) \\=P( Z_\bar{X} \geq 2.142857) }= 0.0160622 $$

jimthompson5910 (jim_thompson5910):

I'm getting 0.0161 roughly, so I agree with perl's answer

OpenStudy (anonymous):

P(X>1200) = P(Z>0.21428571428571427) = 0.4151621288278977 = 1/2.41 ?

jimthompson5910 (jim_thompson5910):

you should have 2.142857 and not 0.2142857

OpenStudy (anonymous):

so my answer is 0.0160622

jimthompson5910 (jim_thompson5910):

approximately, yes

OpenStudy (perl):

1/2.41 is too big ... that is 0.4149

OpenStudy (anonymous):

K...so my answer is 0.0160622

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

Thank you, that's what i will put

OpenStudy (perl):

0.0161 to 4 decimals